The income elasticity of demand of commodity M, elasticity income is 0.48

a Interpret the magnitude in the elasticity value

b.Write the formula for calculating such an elasticity (define all symbols)

c.By how much will demand of commodity M change if income should decrease by 12.5%


1
Expert's answer
2021-05-16T17:43:17-0400

Solution:

a.). The income elasticity of demand measures how responsive the quantity demanded for a good or service is to a change in income. That is, it measures how quantity demanded changes in response to income.

Commodity M has an income elasticity of demand of 0.48, which suggests that it is income inelastic. This means that an increase in income leads to a smaller percentage increase in demand. Consumers are not sensitive to changes in their income when it comes to purchasing commodity M and the demand does not fall significantly with a fall in income. Commodity M is, therefore, a normal good, where 0>YED<1.

 

b.). The formula for calculating income elasticity of demand is the percentage change in quantity demanded divided by the percentage change in income.

Income elasticity of demand (YED) = "\frac{\%\triangle Qd}{\%\triangle Y}"

YED = "=\frac{(Q_{2} -Q_{1})}{(Q_{2}+Q_{1})} \div \frac{(I_{2} -I_{1})}{(I_{2}+I_{1}) }"


Where: Q2 = New quantity demanded

            Q1 = Old quantity demanded

             I2 = New Income

             I1 = Old Income

           YED = Income elasticity of demand.

 

c.). YED ="\frac{\%\triangle Qd}{\%\triangle Y}"


0.48 = "\frac{\%\triangle Qd}{-12.5\%}"


%Δ Qd  = -12.5% "\times" 0.48 = -6%

%Δ Qd = -6%

The demand for commodity M will decrease by 6% when income decrease by 12.5%